OpenStudy (anonymous):

The length and width of a rectangle are 17 cm and 9 cm. If the dimensions are tripled, by exactly how much has the perimeter of the rectangle increased? (Hint: Use the Numeric Method) 26 cm 104 cm 78 cm 52 cm

OpenStudy (anonymous):

help

OpenStudy (anonymous):

its 78 (:

OpenStudy (anonymous):

help please show me how

OpenStudy (ajprincess):

The length of the new rectangle=17*3 =51 The width of the new rectangle=9*3 =27 Perimeter=2(length+width) =2(17+9) =2(26) Perimeter of the new rectangle=2(length+width) =2(51+27) =2(78) The perimeter is increased by 2(78)-2(26). 2(78)-2(26)=2(78-26) =2(52) =104

OpenStudy (anonymous):

shes correct i forgot about the last part